OpenStudy (anonymous):

Verify the identity ((1-sin x)/cos x) = (cosx/(1+sin x))

hero (hero):

\[ (1-\sin x)(1 + \sin x) = \cos^2x \\1 - \sin^2x = \cos^2x \\1 = \sin^2x + \cos^2x \\1 = 1 \]